John was consuming 100 balls of Cake (X) and 50 pieces of Fish (Y) per month. The price of Cake rose from GHS 2 to GHS 3. The price of Fish remained the same at GHS 4. How much would John’s income (M) have to rise so that he can still exactly afford 100 cakes and 50 fish?


1
Expert's answer
2021-05-31T14:59:00-0400

​​​​​​INITIALLY

CAKE QUANTITY - 100

​CAKE PRICE - 2 GHS

FISH QUANTITY- 50

FISH PRICE - 4 GHS

BUDGET = 2(100)+4(50)

= 200+200=400GHS

IF CAKE PRICE RISE TO 3GHS ,IT NEEDS 300 GHS TO BUY 100 QUANTITY that makes the budget 300+200-=500GHS for John to buy same quantity of each.

John’s income (M) have to rise so that he can still exactly afford 100 cakes and 50 fish is= required income-present income= 500-400 i.e 100GHS